# The Effect of Personalization in FedProx: A Fine-grained Analysis on Statistical Accuracy and Communication Efficiency

### v2

@@ A.2 Proof of global statistical Error Bound
To facilitate our analysis, we denote
$\bm{\delta}_{\star}^{(i)}:={{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}},\quad\tilde{\bm{\delta}}^{(i)}:={\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}.$ (20)
that is, $\bm{\delta}_{\star}^{(i)}$ measures the difference between the true global model ${{\bm{w}}_{\star}^{(g)}}$ and the true local model ${{\bm{w}}_{\star}^{(i)}}$, and $\tilde{\bm{\delta}}^{(i)}$ is the estimator of such a difference.
To prove the global statistical error bound, we start with the optimality condition of ${\tilde{\bm{w}}^{(i)}}$ and ${\tilde{\bm{w}}^{(g)}}$, which yields
@@
For the first term on the RHS in (22), apply the $\mu$-strongly convexity of the loss function $\ell$ (Assumption 2)
$\begin{split}\sum_{i\in[m]}\sum_{j\in[n_{i}]}\frac{p_{i}}{n_{i}}\left(\ell({\tilde{\bm{w}}^{(i)}},z_{ij})-\ell({{\bm{w}}_{\star}^{(i)}},z_{ij})\right)&\geq-\sum_{i\in[m]}\sum_{j\in[n_{i}]}\frac{p_{i}}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle\\ &+\mu\sum_{i\in[m]}\sum_{j\in[n_{i}]}\frac{p_{i}}{n_{i}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2},\\ \end{split}$ (23)
@@
$\begin{split}0&\geq-\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle+\mu\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}\\ &+\sum_{i\in[m]}p_{i}\frac{\lambda}{2}\|{\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}\|^{2}-\sum_{i\in[m]}p_{i}\frac{\lambda}{2}\|{{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}.\end{split}$ (24)
@@
00 0 $\geq-\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle+\mu\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\sum_{i}p_{i}\left\|\tilde{\bm{\delta}}^{(i)}\right\|^{2}-\frac{\lambda}{2}R^{2}.$ (25)
Utilizing the result (25) in Lemma 1, we are ready to proof Theorem 1
Proof.
Based on the result in (25), we will discuss the upper bound of $\tilde{\Delta}^{(g)}_{\text{stat}}$ with two cases of statistical heterogeneoity:
Case 1: for the first case, if $R$ is large, i.e, $R>\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}$.
First, applying Cauchy inequality on the first term of (25), it yields
@@
$\begin{split}\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\mathbb{E}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle\leq\sum_{i}\rho\frac{p_{i}}{\sqrt{n_{i}}}\|\tilde{\Delta}^{(i)}_{\text{stat}}\|_{2}.\\ \end{split}$ (28)
@@
$0\geq-\rho\sum_{i}\frac{p_{i}}{\sqrt{n_{i}}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}+\mu\sum_{i}p_{i}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\sum_{i}p_{i}\left\|\tilde{\bm{\delta}}^{(i)}\right\|^{2}-\frac{\lambda}{2}R^{2}.$ (29)
@@
$0\geq-\rho\left(\Big{(}\sum_{i\in[m]}\frac{p_{i}}{n_{i}}\Big{)}\Big{(}\sum_{i}p_{i}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}\Big{)}\right)^{\frac{1}{2}}+\mu\sum_{i}p_{i}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}-\frac{\lambda}{2}R^{2}.$ (30)
@@
The equation (30) would be regarded as a quadratic function of $\sum_{i}p_{i}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}$
Set the $\lambda$ as
$\lambda=\frac{\rho^{2}}{R^{2}}\sum_{i\in[m]}\frac{p_{i}}{n_{i}}.$ (31)
@@
Solving the inequality in (30) w.r.t $\sum_{i}p_{i}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}$, it yields
@@
$\begin{split}\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|_{2}&\leq\left(\sum_{i}p_{i}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}\right)^{\frac{1}{2}}\leq\frac{1}{2\mu}\left(\rho\left(\sum_{i\in[m]}\frac{p_{i}}{n_{i}}\right)^{\frac{1}{2}}+\left(\rho^{2}\sum_{i\in[m]}\frac{p_{i}}{n_{i}}(1+2\mu)\right)^{\frac{1}{2}}\right)\\ &=\frac{(1+\sqrt{1+2\mu})\rho}{2\mu}\left(\sum_{i\in[m]}\frac{p_{i}}{n_{i}}\right)^{\frac{1}{2}}.\\ \end{split}$ (32)
Case 2: for the second case, if $R$ is relatively small, which means $R\leq\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}$, based on the result (25), we know
$\begin{split}0&\geq\sum_{i\in[m]}\sum_{j\in[n_{i}]}\frac{p_{i}}{n_{i}}\left[\ell\left({\tilde{\bm{w}}^{(i)}},x_{ij}\right)-\ell({{\bm{w}}_{\star}^{(i)}},x_{ij})\right]+\frac{\lambda}{2}\sum_{i\in[m]}p_{i}\left(\left\|{\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}\right\|^{2}-\left\|{{\bm{w}}_{\star}^{(g)}}-\bm{w}_{\star}^{(i)}\right\|^{2}\right)\\ &\geq-\underbrace{\sum_{i\in[m]}\sum_{j\in[n_{i}]}\frac{p_{i}}{n_{i}}\left\langle\nabla\ell({\tilde{\bm{w}}^{(i)}},x_{ij}),\tilde{\Delta}^{(g)}_{\text{stat}}\right\rangle}_{T_{1}}+\underbrace{\sum_{i\in[m]}\sum_{j\in[n_{i}]}\frac{p_{i}}{n_{i}}\left\langle\nabla\ell({\tilde{\bm{w}}^{(i)}},x_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}-\tilde{\Delta}^{(g)}_{\text{stat}}\right\rangle}_{T_{2}}\\ &\quad+\sum_{i\in[m]}\sum_{j\in[n_{i}]}\frac{p_{i}}{n_{i}}\mu\|\tilde{\Delta}^{(i)}_{\text{stat}}\|^{2}+\frac{\lambda}{2}\sum_{i\in[m]}p_{i}\|\tilde{\bm{\delta}}^{(i)}\|^{2}-\frac{\lambda}{2}\sum_{i\in[m]}p_{i}\|\bm{\delta}_{\star}^{(i)}\|^{2}.\end{split}$ (33)
It is necessary to remark that in the case of $R\leq\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}$, we should take the GlobalTrain (2) into account so that our analysis different from the case 1.
For $T_{1}$
$T_{1}\leq\big{\|}\sum_{ij}\frac{p_{i}}{n_{i}}\nabla\ell({\tilde{\bm{w}}^{(i)}},z_{ij})\big{\|}_{2}\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|_{2}\leq\rho\sqrt{\sum_{i}\frac{p_{i}^{2}}{n_{i}}}\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|_{2},$ (34)
where the first inequality comes from Cauchy inequality and to obtain the second inequality, we need to note the fact that from Assumption 3, we have
$\left(\mathbb{E}\left\|\sum_{ij}\frac{p_{i}}{n_{i}}\nabla\ell(\widetilde{\bm{w}}^{(i)},z_{ij})\right\|_{2}\right)^{2}\leq\mathbb{E}\left\|\sum_{ij}\frac{p_{i}}{n_{i}}\nabla\ell(\widetilde{\bm{w}}^{(i)},z_{ij})\right\|_{2}^{2}\leq\sum_{ij}\frac{p_{i}^{2}}{n_{i}^{2}}\rho^{2}.$ (35)
Analogy to the induction of $T_{1}$, we can induce $T_{2}$ in a similar way
$\begin{split}T_{2}&\leq\sum_{i}p_{i}\left\langle\frac{1}{n_{i}}\sum_{j}\nabla\ell({\tilde{\bm{w}}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}-\tilde{\Delta}^{(g)}_{\text{stat}}\right\rangle\\ &\leq\sum_{i}p_{i}\left\|\frac{1}{n_{i}}\sum_{j}\nabla\ell({\tilde{\bm{w}}^{(i)}},z_{ij})\right\|_{2}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}-\tilde{\Delta}^{(g)}_{\text{stat}}\right\|_{2}\\ &\leq\sum_{i}p_{i}\frac{\rho}{\sqrt{n_{i}}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}-\tilde{\Delta}^{(g)}_{\text{stat}}\right\|_{2}.\\ \end{split}$ (36)
Note the fact
$\left\|\tilde{\Delta}^{(i)}_{\text{stat}}-\tilde{\Delta}^{(g)}_{\text{stat}}\right\|_{2}=\left\|{{\bm{w}}_{\star}^{(i)}}-{\tilde{\bm{w}}^{(i)}}-{{\bm{w}}_{\star}^{(g)}}+{\tilde{\bm{w}}^{(g)}}\right\|_{2}=\left\|\bm{\delta}_{\star}^{(i)}-\tilde{\bm{\delta}}^{(i)}\right\|_{2}.$ (37)
Combining (37) back to (36), it yields
$T_{2}\leq\sum_{i}p_{i}\frac{\rho}{\sqrt{n_{i}}}\left\|\bm{\delta}_{\star}^{(i)}-\tilde{\bm{\delta}}^{(i)}\right\|_{2}\leq\sum_{i}p_{i}\frac{\rho}{\sqrt{n_{i}}}\left(\left\|\bm{\delta}_{\star}^{(i)}\right\|_{2}+\left\|\tilde{\bm{\delta}}^{(i)}\right\|_{2}\right).$ (38)
Plugging (34) and (38) into (33), it yields
$\begin{split}0&\geq-\rho\sqrt{\sum_{i}\frac{p_{i}^{2}}{n_{i}}}\|\tilde{\Delta}^{(g)}_{\text{stat}}\|_{2}-\sum_{i}p_{i}\frac{\rho}{\sqrt{n_{i}}}(\|\bm{\delta}_{\star}^{(i)}\|_{2}+\|\tilde{\bm{\delta}}^{(i)}\|_{2})+\mu\sum_{i}p_{i}\|\tilde{\Delta}^{(i)}_{\text{stat}}\|^{2}\\ &\quad\quad+\frac{\lambda}{2}\sum_{i}p_{i}\|\tilde{\bm{\delta}}^{(i)}\|^{2}-\frac{\lambda}{2}\sum_{i\in[m]}p_{i}\|\bm{\delta}_{\star}^{(i)}\|^{2}\\ &\geq\mu\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}-\rho\sqrt{\sum_{i}\frac{p_{i}^{2}}{n_{i}}}\|\tilde{\Delta}^{(g)}_{\text{stat}}\|_{2}-\rho\sum_{i}\frac{p_{i}}{\sqrt{n_{i}}}\|\tilde{\bm{\delta}}^{(i)}\|_{2}+\frac{\lambda}{2}\sum_{i}p_{i}\|\tilde{\bm{\delta}}^{(i)}\|^{2}-R\rho\sum_{i}\frac{p_{i}}{\sqrt{n_{i}}}-\frac{\lambda}{2}R^{2},\end{split}$ (39)
where the last line uses Assumption 1. Then focusing on the right-hand side of the above inequality and regarding it as a function of $\tilde{\bm{\delta}}^{(i)}$, it can be written as

### v3

@@ A.3.2 Proof of Global Statistical Error Bound in Theorem 1
We start with the optimality condition of ${\tilde{\bm{w}}^{(i)}}$ and ${\tilde{\bm{w}}^{(g)}}$, which yields
@@
For the first term on the RHS in (25), apply the $\mu$-strongly convexity of the loss function $\ell$ (cf. Assumption 3), we then have
$\begin{split}\sum_{i\in[m]}\sum_{j\in[n_{i}]}\frac{p_{i}}{n_{i}}\left(\ell({\tilde{\bm{w}}^{(i)}},z_{ij})-\ell({{\bm{w}}_{\star}^{(i)}},z_{ij})\right)&\geq-\sum_{i\in[m]}\sum_{j\in[n_{i}]}\frac{p_{i}}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle\\ &+\frac{\mu}{2}\sum_{i\in[m]}\sum_{j\in[n_{i}]}\frac{p_{i}}{n_{i}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2},\\ \end{split}$ (26)
@@
$\begin{split}0&\geq-\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle+\frac{\mu}{2}\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}\\ &+\sum_{i\in[m]}p_{i}\frac{\lambda}{2}\|{\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}\|^{2}-\sum_{i\in[m]}p_{i}\frac{\lambda}{2}\|{{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}.\end{split}$ (27)
@@
00 0 $\geq-\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle+\frac{\mu}{2}\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\sum_{i}p_{i}\left\|\tilde{\bm{\delta}}^{(i)}\right\|^{2}-\frac{\lambda}{2}R^{2}.$ (28)
Applying Cauchy inequality on the first term of (28), it yields
@@
$\begin{split}\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\mathbb{E}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle\leq\sum_{i}\rho\frac{p_{i}}{\sqrt{n_{i}}}\mathbb{E}\|\tilde{\Delta}^{(i)}_{\text{stat}}\|_{2}.\\ \end{split}$ (31)
@@
$0\geq-\rho\sum_{i}\frac{p_{i}}{\sqrt{n_{i}}}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}+\frac{\mu}{2}\sum_{i}p_{i}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\sum_{i}p_{i}\mathbb{E}\left\|\tilde{\bm{\delta}}^{(i)}\right\|^{2}-\frac{\lambda}{2}R^{2}.$ (32)
@@
$0\geq-\rho\left(\Big{(}\sum_{i\in[m]}\frac{p_{i}}{n_{i}}\Big{)}\Big{(}\sum_{i}p_{i}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}\Big{)}\right)^{\frac{1}{2}}+\frac{\mu}{2}\sum_{i}p_{i}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}-\frac{\lambda}{2}R^{2}.$ (33)
@@
The equation (33) would be regarded as a quadratic function of $\sum_{i}p_{i}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}$. If we set the $\lambda$ as
$\lambda\leq\frac{\rho^{2}}{\mu R^{2}}\sum_{i\in[m]}\frac{p_{i}}{n_{i}},$ (34)
@@
then solving the inequality in (33) w.r.t $\sum_{i}p_{i}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}$ yields
@@
$\begin{split}\mathbb{E}\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|_{2}&\leq\left(\sum_{i}p_{i}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}\right)^{\frac{1}{2}}\leq\frac{1}{\mu}\left(\rho\left(\sum_{i\in[m]}\frac{p_{i}}{n_{i}}\right)^{\frac{1}{2}}+\left(\rho^{2}\sum_{i\in[m]}\frac{3p_{i}}{n_{i}}\right)^{\frac{1}{2}}\right)\\ &=\frac{(1+\sqrt{3})\rho}{\mu}\left(\sum_{i\in[m]}\frac{p_{i}}{n_{i}}\right)^{\frac{1}{2}}.\\ \end{split}$ (35)
Case 2: Then we consider $R\leq\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}$. Recall the definition of the global minimizer of GlobalTraining $\tilde{\bm{w}}_{\text{GT}}=\arg\min_{w}\sum_{i}p_{i}L_{i}(w;S_{i})$, based on the optimality condition of $\widetilde{\bm{w}}_{i},\widetilde{\bm{w}}_{g}$, $i\in[m]$, we obtain